on a newly discovered planet, the period of a pendulum with a length of 2 m is found to be 5 seconds. what is the acceleration due to gravity on this new planet?
Answer #1

T = 5s

L = 2m

Since T = 2.pai √L/g

=) (T/2π )^2 = L/g

=) g = L ( 2π/T)^2 = 2×4×3.143×3.143/25 = 3. 16m/s^2

=) g = 3.16m/s^2

Plz give a thumb up. Thank you.